Phobos, the larger of Mars’ two moons, orbits 5.99×10 3 km above the planet surface.
Given that the mass of Mars is 6.42×10 23 kg and the radius of Mars is 3.39×10 3 km, what are
the orbital speed and period of Phobos?
1
Expert's answer
2020-05-12T09:57:02-0400

Let denote the height at which Phobos is revolving from the surface of the Mars "h" .Radius of Mars is "R", mass of Mars and Phobos are "M \&m" respectively,"v" is the orbital velocity of Phobos and "T" is the orbital period of Phobos.

We have given that,

"h=5.99\times 10^3km=5.99\times 10^6m\\
M=6.42\times10^{23}kg\\
R=3.39\times 10^3km=3.39\times 10^6m"

Now, we know that from Newton's law of Gravitation, The gravitational force(centripetal force)of Mars is the only force which responsible for revolution of Phobos around Mars,thus

Force between Mars and Phobos is

"F=G\frac{Mm}{(R+h)^2} \hspace{1cm} (1)"

where, "R+h" is the distance from center of Mars to Phobos.hence, the same force "F" can also be written as

"F=\frac{mv^2}{(R+h)} \hspace{1cm}(2)"

Therefore, from equation 1 and 2 we get,


"G\frac{Mm}{(R+h)^2}=\frac{mv^2}{(R+h)}\\
\implies v=\sqrt{\frac{GM}{R+h}} \hspace{1cm} (\clubs)"

Now, from equation "(\clubs)" ,


"v=\sqrt{\frac{6.67\times 10^{-11}\cdot 6.42\times 10^{23}}{9.38 \times 10^6}}\\
v=2.136\times 10^3ms^{-1}"

Therefore, orbital speed of Phobos is "v=2.136\times 10^3ms^{-1}" .


Now, we also know that

"v=(R+h)\omega\\
\implies v= (R+h) \frac{2\pi}{T}\\
\implies T=\frac{2\pi (R+h)}{v} \hspace{1cm}(\spades)"

Thus, from equation "(\spades)" we get,


"T=\frac{2\pi \times 9.38\times 10^6}{2.136\times10^3}\\
\implies T=27.591\times10^3 s"

Therefore, period of Phobos is "T=27.591\times10^3 s" .